### v4

# OpsEval: A Comprehensive Benchmark Suite for Evaluating Large Language Models’ Capability in IT Operations Domain

Abstract
The past decades have witnessed the rapid development of Information Technology (IT) systems, such as cloud computing, 5G networks, and financial information systems. Ensuring the stability of these IT systems has become an important issue. Large language models (LLMs) that have exhibited remarkable capabilities in NLP-related tasks are showing great potential in AIOps, such as root cause analysis of failures, generation of operations and maintenance scripts, and summarizing of alert information. Unlike knowledge in general corpora, knowledge of Ops varies with the different IT systems, encompassing various private sub-domain knowledge, sensitive to prompt engineering due to various sub-domains, and containing numerous terminologies. As a consequence, existing NLP-related benchmarks such as C-Eval and MMLU can not guide the selection of suitable LLMs for Ops, and current metrics like BLEU and ROUGE can not adequately reflect the question-answering (QA) effectiveness in the Ops domain. Therefore, this paper proposes a comprehensive benchmark suite named OpsEval, including an Ops-oriented evaluation dataset, an evaluation benchmark for Ops, and a specially designed QA evaluation method for Ops. Our dataset contains 7,184 multiple-choice questions and 1,736 QA questions. We have carefully selected and released 20% of the dataset by CC-BY-NC-4.0 license [^1] Data page is available at https://github.com/NetManAIOps/OpsEval-Datasets. written by domain experts in various sub-domains to assist current researchers in preliminary evaluations of their LLMs tailored for Ops (OpsLLM). The remaining undisclosed 80% of the data is used to prevent test set leakage. We test over 21 latest LLMs under various settings such as self-consistency, chain-of-thought, and in-context learning, revealing findings when applying LLMs to Ops. We also propose an evaluation method for QA in Ops, which has a coefficient of 0.9175 with human experts and is improved by 0.2470 and 1.313 compared to BLEU and ROUGE, respectively. Over the past seven months, our dataset and leaderboard [^2] Leaderboard is available at https://opseval.cstcloud.cn/content/leaderboard. have been continuously updated. For reproducibility, the evaluation framework code is publicly available.
1 Introduction
The IT Operations (Ops) plays a crucial role in maintaining the efficient and stable operation of information systems such as cloud computing, 5G networks [^3] Strictly speaking, 5G belongs to the field of communications technology (CT), but given its broad association with the information technology (IT) sector, for the sake of generality, we refer to it as IT operations, abbreviated as Ops, throughout the remainder of this paper. and financial information systems. As the Internet continues to expand rapidly, the scale and complexity of systems are escalating, leading to the emergence of artificial intelligence-assisted operations as a novel trend. Termed “AIOps” by Gartner (Lerner, 2017), this technique utilizes artificial intelligence to address (but is not limited to) tasks such as anomaly detection, fault analysis, generation of alert summaries, performance optimization, and capacity planning.
In recent years, large language models (LLMs) have witnessed significant advancements. The latest models, such as GPT-4o (OpenAI, 2024), GPT-4V (OpenAI, 2023b), Meta-Llama-3 (AI@Meta, 2024), and GLM-4 (Zeng et al., 2022), have demonstrated exceptional generalization and task-planning capabilities. As a result, these models have provided numerous opportunities to enhance downstream domain-specific applications. With its advanced summarizing, report analyzing, and ability to diagnose errors, LLM is well suited for Ops on tasks like question answering, information summarizing, and report analysis. Hereinafter, we refer to the LLM used for Ops as OpsLLM, regardless of whether they have been optimized specifically for Ops.
While there are benchmarks for assessing general-purpose NLP-related capabilities, such as C-EVAL (Huang et al., 2023) and AGIEval (Zhong et al., 2023), there are also benchmarks for specific domains, like FinEval (Zhang et al., 2023) in the financial sector and CMB (Wang et al., 2023a) in the medical sector. However, no benchmark exists to evaluate the effectiveness of LLMs or OpsLLMs in Ops tasks. There is an urgent need for an Ops benchmark that informs us about the performance of current LLMs on Ops tasks. On the other hand, a good benchmark can significantly aid the optimization process of OpsLLMs tailored for the Ops domain. Nevertheless, due to the specialty of the Ops tasks, constructing an Ops benchmark presents the following challenges:
  • Sensitive data. The Ops data is primarily sensitive and proprietary to companies, with very few publicly available data, making it difficult for any company to independently provide sufficient evaluation data to ensure confidence in the test results.
  • Sub-domains. There are many different sub-domains, such as 5G communications, cloud computing, and bank transaction systems. Ops within each sub-domain typically requires a combination of various capabilities, such as network configuration or terminology explanation, which we refer to as “tasks”. Due to the large number of sub-domains and tasks in the Ops field and the lack of an authoritative and systematic taxonomy, classifying a large number of questions becomes a challenge.
  • Prompt sensitivity. Due to the relatively proprietary nature of the Ops, existing LLMs have not undergone specialized supervised fine-tuning (SFT) for instruct following within the Ops field, the evaluation results are more sensitive to prompt engineering. Designing appropriate prompts for robust and accurate evaluation is challenging.
  • QA metric. Existing metrics like BLEU only consider the similarity of model output to the reference in natural language aspects, which does not always reflect true performance in Ops tasks. Some terms and expressions in Ops sub-domains have specific meanings that LLM cannot summarize. Designing an automatic evaluation method that assesses the accuracy of QA of Ops from an accurate semantic level is challenging.
To address these issues, we propose OpsEval, a comprehensive benchmark suite for evaluating LLMs’ capability in the IT operations domain. First, to tackle the challenge of benchmark data mostly being private and not publicly shareable, we initiated a community around AIOps, which has attracted dozens of companies to participate. We have selected 10 representative sub-domains from the community, allowing continuous data contributions from community members. We then aggregate data under the same sub-domain to ensure robustness in evaluation. Additionally, we generated multi-choice (MC) and question-answering (QA) questions as supplements based on publicly available network management books. To address the challenge of classifying the numerous sub-domains and tasks in the Ops field, we employ model-based pre-clustering and manual review to annotate eight tasks and three abilities for independent evaluation. Considering the prompt sensitivity of benchmark results, we systematically test model performance under self-consistency (SC), chain-of-thought (CoT), and few-shot in-context learning (ICL). The prompts used in our evaluation are also disclosed in the paper. Lastly, to address the inaccuracy of existing metrics in Ops QA evaluation, we design FAE-Score, which assesses questions from the perspectives of fluency, accuracy, and evidence. Experimental results show that our designed metrics align strongly with the annotations of human experts.
The contributions of our paper are as follows: 1) We introduce OpsEval, the first bilingual multi-task dataset in the operations domain, covering 8 tasks and 3 abilities with 8920 questions. To assist researchers in preliminary evaluating their OpsLLMs, we have carefully selected and released 20% of QAs from our benchmark licensed under CC-BY-NC-4.0, with the remaining 80% of undisclosed data preventing unfair evaluations due to data leakage (Wei, et.al., 2023) 2) Based on the dataset, we introduce the OpsEval evaluation benchmark, conducting independent and robust evaluations with various prompting techniques and a specifically designed evaluation metric named FAE-Score. Compared to the commonly employed BLEU and ROUGE metrics, FAE-Score exhibits a more pronounced congruence with the evaluations of human experts. Specifically, FAE-Score attains a correlation coefficient 0.9175 with expert assessments, surpassing the coefficients of 0.6705 for BLEU and -0.3957 for ROUGE. 3) We released an online leaderboard that continuously updates mainstream LLMs’ performance on Ops tasks for the past 7 months. To ensure the reproducibility and reliability of this leaderboard, we made our evaluation framework’s code publicly available [^4] Evaluation framework code is available at https://github.com/NetManAIOps/OpenCompass-OpsQA.. 4) Based on the results of OpsEval evaluation, we provide key observations and practical lessons to help domain practitioners make decisions such as whether existing models are sufficiently applicable within a specific sub-domain, the necessity for fine-tuning and whether model quantization compromises the effectiveness.
2 Related Works
As LLMs evolve rapidly, their complex and varied capabilities are increasingly recognized. As a result, there is a growing trend towards evaluation benchmarks tailored specifically for LLMs. These can be divided into two categories: general ability benchmarks and domain-specific benchmarks.
Table 1: A comparison of OpsEval with other popular datasets/benchmarks.
MMLU HELM BIG-bench SEAL C-Eval AGIEval FLUE MultiMedQA CMB NetOps OpsEval
Ops Domain Dataset
Open-sourced Benchmark
Up-to-date Leaderboard
General ability benchmarks assess the general abilities of LLMs across various tasks. These tasks evaluate LLMs’ capacity for logical reasoning, general knowledge, common sense, and other similar abilities rather than being confined to a particular domain. MMLU (Hendrycks et al., 2021) is a benchmark designed to measure knowledge acquired during pretraining by evaluating models exclusively in zero-shot and few-shot settings, covering 57 subjects across STEM. HELM (Liang et al., 2022) employs seven distinct metrics in 42 unique scenarios, offering a comprehensive evaluation of LLMs’ capabilities across multiple dimensions. BIG-bench (Srivastava et al., 2022) comprises 204 tasks spanning a wide array of topics, with a particular focus on tasks deemed beyond the reach of current LLMs. SEAL (AI, 2024b) features private, expert evaluations of leading frontiers models. C-Eval (Huang et al., 2023) is a comprehensive Chinese evaluation suite designed to assess Chinese LLMs’ advanced knowledge and reasoning abilities rigorously. AGIEval (Zhong et al., 2023) curates authentic questions from examinations such as the Chinese College Entrance Exam (CCEE) and the SAT, constructing a fundamentally human-centric evaluation dataset.
Domain-specific benchmarks evaluate the abilities of LLMs to handle tasks in specific fields. These benchmarks require LLMs to possess specialized knowledge in a specific domain and to respond in a manner consistent with the cognitive patterns of that field. Despite the rapid progression of LLMs in specialized domains, the evaluation metrics for these specific areas have received less attention. FLUE (Shah et al., 2022) is an open-source comprehensive suite of benchmarks, including new benchmarks across 5 NLP tasks in financial domain. MultiMedQA (Singhal et al., 2022) is an extensive medical question-answering dataset, with questions derived from professional medical exams, research, and consultation records. CMB (Wang et al., 2023a) includes multi-choice questions (CMB-Exam) and complex clinical questions based on real case studies (CMB-Clin). NetOps (Miao et al., 2023) focuses on evaluations in the network field, which is relevant to the field of Ops. NetOps includes multi-choice questions in both English and Chinese and a few cloze tests and question-answering questions. However, they only focus on wired network operations and while the dataset is released, they lack a benchmark that continuously updates the leaderboard.
3 OpsEval Benchmark
Figure 1: The framework of OpsEval
Figure 1 shows the overall framework of OpsEval from construction to evaluation. We collected data from multiple sources and then preprocessed it to enhance its quality. Finally, we evaluated LLMs on the dataset using various prompt engineering techniques.
3.1 Data Collection
Our benchmark questions have been collected from various sources; we summarize them into four categories: company materials, certification exams, Ops textbooks, and automated generation. Each source is highly esteemed globally and reviewed by our Ops collaborators.
Company Materials. include production environment materials like Ops tickets and error logs , as well as internal documents and tests for Ops staff training. We have established cooperative relationships with 10 companies, covering various sectors like telecommunications, finance, and Ops service/tool providers, and received expert collaboration and Ops materials from them. The Appendix A.1 provides information about the companies and experts.
Certification Exams. include knowledge assessments necessary for becoming an Ops staff and are naturally in the form of multiple-choice and question-answering questions. We obtained the relevant study guidebooks for these certification exams from public book websites and extracted sample questions from them as one of the sources for Ops questions.
Operations Textbooks. We first constructed a seeding keyword list for the Ops field and searched for related books. The textbooks contain relatively complete knowledge content, which can provide experts with materials for question creation, and some books themselves also include a certain number of exercises at the end of the chapters.
Automated Generation. To enhance the diversity and depth of our test set, we source QAs from authoritative books covering a range of Ops domains by extracting textbook contents and asking GPT4 to generate questions. In Appendix A.7, we discussed the methodology and challenges.
Table 2: The number of questions in OpsEval, grouped by their sub-domains.
Sub-domain Source License Type Questions
Wired Network Operation Textbooks Public Multi-Choice 3901
5G Communication Certification Exams Public Multi-Choice 2615
Question-Answering 1162
Oracle Database Company Materials CC-BY-NC-4.0 Multi-Choice 497
Log Analysis Company Materials CC-BY-NC-4.0 Question-Answering 420
DevOps Automated Generation CC-BY-NC-4.0 Question-Answering 154
Securities Information System Company Materials CC-BY-NC-4.0 Multi-Choice 91
Hybrid Cloud Company Materials CC-BY-NC-4.0 Multi-Choice 40
Financial IT Company Materials CC-BY-NC-4.0 Multi-Choice 40
3.2 Preprocessing
We systematically carried out the preprocessing of our original data in the following stages:
Deduplication: Any repeated or highly similar questions are identified and removed to avoid redundancy in the test set. We calculate the cosine similarity of the question stems by bge-large-zh-v1.5  (Xiao et al., 2023) to detect duplicate questions and identify pairs of questions with a similarity above a certain threshold (th=0.7).
Dependance Filtering: We have filtered out questions that rely on external images or document content to ensure the completeness of the question content itself. The filtering process was done by two parallel lists of empirical keywords in the question stems and the responses of GPT-3.5-turbo. The keyword list can be found in the Appendix A.2.
Question Categorization: We devise a categorization that captures many tasks that professionals confront in practical applications. The categorization process consists of two steps: automated screening and manual review. We first use GPT-4 for topic modeling to gain rough insights about the dataset and determine the relevance of each question to Ops, which resulted in more than 20 tasks but had an imbalanced distribution [^5] For prompts used during the preprocessing, please see in Appendix.. We then involved dozens of experts during the manual review process to categorize the questions into eight tasks and three abilities. The distribution of the questions across these eight tasks and three ability levels is shown in Table 3, and the details of each task and ability can be found in Appendix A.3.
Table 3: The distribution of different tasks and abilities of questions in OpsEval.
Category Percentage (%)
Task Automation Scripts $3.3$
Monitoring and Alerting $5.2$
Performance Optimization $5.3$
Software Deployment $7.9$
Fault Analysis and Diagnostics $13.7$
Network Configuration $29.0$
General Ops Knowledge $20.2$
Miscellaneous $15.5$
Ability Knowledge Recall $49.8$
Analytical Thinking $39.9$
Practical Application $10.2$
Manual Review: In the manual review step, we asked Ops experts from the industry to inspect the results of the previous three automated steps, including confirming duplicate and invalid questions and examining the classification results of GPT-4. In our work, an expert is defined as an individual with ten or more years of professional experience in their field, whether as an employee or a researcher. Experts were also asked to drop the questions unrelated to the Ops field. We split the dataset by n-folds and ensure each fold has at least two experts to review. As listed in Table 2, this quality enhancement process resulted in a refined test set of approximately 7,000 multi-choice and 2,000 question-answering questions.
3.3 Evaluation Settings
Multi-choice questions offer a structured approach with definitive answers. These questions are straightforward and provide a clear metric for assessment. We use accuracy as the metric. A choice-extracting function based on regular expressions is used to extract the predicted answer of LLMs. Then, we calculate the accuracy based on the extracted answer and the ground-truth labels.
Question-answering questions do not come with predefined options. We use two metrics for question-answering questions: one is based on word overlaps, and the other is based on semantic similarity. For the first type, we use ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002). For the second type, we use LLM judges and human experts to evaluate the outputs of LLMs, called FAE-Score and Expert-Evaluation, designed explicitly in OpsEval. For these two metrics, we design three criteria highly related to Ops’ needs. The three criteria in consideration are as follows:
  • Fluency. Assessment of the linguistic fluency in the model’s output and compliance with the question-answering question’s answering requirements.
  • Accuracy. Evaluation of the precision and correctness of the model’s output, including whether it adequately covers key points of the ground-truth answer.
  • Evidence. Examine whether the model’s output contains sufficient argumentation and evidential support to ensure the credibility and reliability of the answer.
For FAE-Score, we provide the judge model with the question and the reference answer, followed by one anonymous model’s output. We use GPT-4 as the judge model of FAE-Score based on evidence that GPT-4 can reliably match human preferences with over 80% agreement (Zheng et al., 2023). In the subsequent validation section §5, we also verified the consistency between GPT-4’s scores and expert scores. Figure 2(b) shows the prompt we used to ask the judge model to score. For Expert-Evaluation, we asked experts to score it between 0 and 3 for each criterion. During the scoring, the raw question, the detailed answer and its key points, and the output of an anonymous model are given at each iteration. For the demographical information of the experts, see Appendix A.1.
((a)) The prompt for GPT-4 initial categorization
((b)) The prompt for GPT-4 Score
Figure 2: Prompts for ChatGPT used in the framework of OpsEval.
Prompting Techniques. We use various settings to evaluate LLMs on OpsEval to get a comprehensive overview of their performance. We evaluate LLMs in zero and few-shot (3-shot) settings. For each setting, we evaluate LLMs in four sub-settings of prompt engineering, that is, naive answers (Naive), self-consistency (SC) (Wang et al., 2023b), chain-of-thought (CoT) (Wei et al., 2023), self-consistency with chain-of-thought (CoT+SC). We set the number of queries in SC to 5.
Models. We evaluate popular LLMs covering different weights from different organizations. The model selection was guided by specific criteria: We aimed to include the latest and most advanced large language models, with a particular focus on those capable of handling Chinese input. The detailed information of all 21 LLMs can be found in Table 9 in Appendix B.1.
Fairness Consideration We are currently making 20% of the data available to the public for Ops community contribution and research purposes, yet for fairness of the evaluation, the complete version of the OpsEval dataset is kept undisclosed. To evaluate a new model, users can submit a Docker image with an initialization script when starting a container based on it. We will run the evaluation automatically and obtain the result on the OpsEval website. Users can choose to disclose their results on the leaderboard of OpsEval or not based on their preference.
This released 20% dataset serves as sample questions so that the researchers will know the types and topics of questions and answers expected by the benchmarks. This allow researchers to gain intuition and insights on how to improve their models. Model developers can use the 20% dataset for locally evaluating by themselves their model’s performance, enabling faster iterations of their model training. The 20% dataset can be used as a seed to generate QA pairs using automatic QA generation algorithms (Wang et al., 2023c), ultimately providing more Ops text data for improving new models.
API Exposure. For large language models, user requests constitute private data that should be carefully protected. We believe that ensuring data privacy and not using such data for subsequent training is a fundamental requirement that all publicly accessible models must meet. The APIs we utilized for model evaluation explicitly guarantee that user data will not be used for model training (Microsoft, 2023; BigModel, 2023; Cloud, 2023). When evaluating closed-source models such as ChatGPT and GLM, the questions in this non-disclosed questions were sent (thus exposed) to their APIs, but the answers to these questions were still non-disclosed.
4 Result Analysis
4.1 Overall Performance
Figure 3: LLMs’ overall performance on Wired Network Operations English test set (3-shot). Models are ranked based on their mean accuracy among different settings. The error bars represent the variance in the model’s accuracy across different prompting techniques.
Table 4: LLMs’ overall performance (Accuracy%) on Wired Network Operations English test set (3-shot). Models are ranked based on their best performance among different settings.
Model Naive SC CoT CoT+SC Best Score Variance
GPT-4 / / 88.70 / 88.70 /
Claude-3-Opus 79.70 78.36 81.24 82.28 82.28 2.9638
Qwen2-72B-Instruct 76.91 76.91 80.12 80.65 80.65 4.0720
Qwen2-7B-Instruct 65.74 65.74 67.42 67.42 67.42 0.9408
GLM-4 64.77 64.77 77.06 77.06 77.06 50.3480
GPT-3.5-turbo 68.30 68.30 70.90 72.50 72.50 4.2800
Qwen-72B-Chat 70.32 70.32 70.13 70.22 70.32 0.0083
ERNIE-Bot-4.0 60.00 60.00 70.00 70.00 70.00 33.3333
LLaMA-2-70B 55.00 56.20 66.80 67.20 67.20 43.5866
Qwen1.5-14B-Chat 52.23 53.52 59.53 64.17 64.17 23.0368
DevOps-Model-14B-Chat 63.85 61.96 41.15 44.01 63.85 139.6604
GLM-3-turbo 59.53 59.53 63.65 63.65 63.65 5.6581
Qwen-14B-Chat 62.60 59.70 50.58 55.88 62.60 26.9921
Meta-LLaMA-3-8B 41.03 42.07 62.45 62.62 62.62 110.2314
LLaMA-2-13B 53.30 53.00 56.80 61.00 61.00 13.9758
InternLM2-Chat-20B 60.48 60.48 45.10 45.10 60.48 78.8481
LLaMA-2-7B 48.20 46.80 52.00 55.20 55.20 14.4366
Qwen-7B-Chat 52.10 51.00 48.30 49.80 52.10 2.6600
Baichuan2-13B-Chat 51.90 51.60 44.50 47.45 51.90 12.5822
Gemma-7B 30.24 30.24 51.56 51.56 51.56 151.5141
InternLM2-Chat-7B 48.2 48.2 49.74 49.74 49.74 0.7905
Mistral-7B 47.22 47.22 45.58 45.58 47.22 0.8965
ChatGLM3-6B 42.10 42.10 43.47 43.47 43.47 0.6256
Gemma-2B 26.63 26.63 37.54 37.54 37.54 39.6760
The results of the few-shot evaluation with four settings on the Wired Network Operation test set are shown in Figure 3. Results of the other sub-domains and settings are shown in Appendix B.4.  [^6] Due to the consideration of time, cost, and API rate limits, for GPT-4, we only make the 3-shot evaluation with the CoT setting to serve as an upper bound of all LLMs to provide a reference. Overall, open-sourced LLMs yield evaluation results on the OpsEval benchmark generally worse than those in general domains like MMLU (Hendrycks et al., 2021) and CEval (Huang et al., 2023). This comparison highlights the necessity of explicitly fine-tuning OpsLLM for the Ops field. Closed source models like GPT-4 and GLM-4 consistently outperform open source models, while smaller models, such as Qwen1.5-14B-Chat, exhibit competitive performance in multi-choice questions, thanks to their fine-tuning process and the quality of their training data. However, their large variance across the four settings suggests that such models may have worse robustness under different prompts. Furthermore, we observed significant variability in how different LLMs respond to various prompt engineering techniques. Given the critical importance of stability in the Ops domain, it is essential to consider a model’s sensitivity to prompts when selecting foundation model. Further research into prompt engineering is needed to improve model performance and reliability in this domain.
Observations: 1) Few-shot and CoT can significantly increase performance if the model is tuned to adapt to these techniques, while SC may have little influence on highly consistent LLMs. 2) Smaller models with weaker natural language abilities are less stable with advanced prompts. Simpler prompts work better for them.
Pratical Lesson: The choice of fundamental models should be a balance between their performance (average score) and robustness (variance) under different prompt settings.
4.2 Performance on Different Tasks and Abilities
Figure 4: LLMs’ performance on eight Ops sub-domains, eight tasks and three abilities. Each colored area presents the lower and upper bound of the corresponding parameter-size group.
Figure 5: Heatmap of failure case distribution regarding tasks and abilities. The values represent the proportion of failure cases across all LLMs; redder areas indicate higher failure rates.
To investigate how LLMs perform in each Ops sub-domain and each task, and to what extent they possess the general abilities, we summarize the result of different parameter-size groups of LLM and plot them on three radar charts in Figure 4. Regarding the eight tasks we tested, LLMs yield higher accuracy in General Knowledge tasks, while their performance drops and varies drastically in highly specialized tasks like Automation Scripts and Network Configuration, reflecting the impact of specialized corpus and domain knowledge on the performance of LLMs. By grouping LLMs by their parameter size, we find that although LLMs with 10B-30B parameters have higher accuracy in their best cases compared with LLMs with no more than 10B parameters, different 10B-20B LLMs’ performance varies drastically. To provide systematic practical lessons for researchers in the operations domain on pre-training and fine-tuning OpsLLM, we have analyzed the error rates of LLMs across the 8 tasks and 3 abilities in Figure 5. By examining the focus areas across different categories, we have identified key research targets for capability training.
Observations: Among the 24 categories of results, models performed the worst in Analytical Thinking for Automation Scripts. This indicates that current models can only recall the learned scripts but struggle to infer their logical relationships. Similarly, Analytical Thinking showed the lowest performance across the three major tasks, indicating that current OpsLLM models still have some way to go before becoming foundational models for Ops Agents. Thus, researchers should focus on inference-related SFT (supervised fine-tuning) datasets.
Insights: 1) Among different sub-domains of Ops, 5G communication and database demand further pretraining and fine-tuning of LLM. 2) To be capable of an Ops agent, the foundation model must be able to make a connection between specialized domain knowledge.
4.3 Performance on Question-Answering
Table 5: LLMs’ performance on English network operations question-answering problems. Total is the sum of the previous three columns.
Model ROUGE BLEU FAE-Score Expert Evaluation
Fluency Accuracy Evidence Total
GPT-3.5-turbo 12.26 6.78 8.47 3.00 1.96 1.20 6.16
LLaMA2-70B 7.74 4.20 7.28 2.92 1.48 1.32 5.72
LLaMA2-13B-Chat 4.98 3.43 7.16 2.82 1.34 1.62 5.78
Baichuan2-13B-Chat 4.76 0.35 5.85 2.40 1.12 1.02 4.54
Qwen-7B-Chat 11.82 4.33 5.63 2.56 1.14 0.84 4.54
ChatGLM3-6B 9.71 5.07 4.88 2.84 0.76 0.76 4.36
InternLM2-7B-Chat 13.27 0.54 4.52 1.80 0.70 0.10 2.60
Table 5 presents the evaluation results of 200 question-answering English questions across four metrics: ROUGE, BLEU, FAE-Score, and Expert-Evaluation, sorted by FAE-Score results. To gain more insight into how different metrics perform in QA evaluation, we use Figure 16 (see in Appendix B.8.2) as a case analysis. While BLEU and ROUGE are efficient in natural language comparison, they lack semantic information to determine which part of the context is more important than others. Knowing that a given benchmark evaluates QA based on BLEU/ROUGE, there is an obvious way to trick the metric: repeat patterns occurring in the question, gaining a higher possibility to match some patterns in the reference answer. Due to their lack of semantic information related to Ops and the potential hack, traditional metrics like BLEU are unsuitable for specialized benchmarks. Instead, with specialized prompting, FAE-Score can pick up those important keywords and not be influenced by repeated words that contain no useful information. In a later section, we discuss the alignment between different metrics and expert evaluation, validating the effectiveness of FAE-Score in automated QA evaluation within the Ops domain.
Practical Lesson: FAE-Score is suitable for large-scale qualitative evaluations in the Ops field.
4.4 Performance on Different Quantization parameters
We conducted experiments on different quantized versions of LLaMA-2-70B and obtained various results and conclusions [^7] For detailed results, please see Appendix B.5.. Overall, although the performance of the INT4 version decreases in both English and Chinese, the decline does not exceed 10%. However, the performance drop in the INT3 version is more significant, requiring careful consideration in practical applications.
Practical Lesson: Quantization with more than 3 bits can effectively reduce computation and memory costs while preserving performance.
5 Validation
5.1 Benchmark Leakage Test
Table 6: Measurement of potential test data leakage during the training of LLM. This demonstrates the unbiased nature and non-leakage of the OpsEval test set.
Dataset $L_{test}$ $L_{ref}$ $\Delta L$
Alpaca 1.994033 2.354260 -0.360228
Alpaca-GPT4 1.498862 1.763663 -0.391062
CEval 2.570809 2.309943 0.260866
MMLU 2.547598 2.189870 0.357728
OpsEval 1.885437 1.728079 0.105095
For the fairness of a benchmark suited for LLM, avoiding potential bias emerging from test set leakage is necessary. We adapted the methodology from Wei, et.al. (2023) to perform a leakage test on OpsEval’s dataset. We evaluate the LLM loss on samples from different datasets for several LLMs and calculate the average loss. For each dataset, we compare LLM loss on the test split ($L_{test}$) and a specially curated reference set ($L_{ref}$ ) generated by GPT-4, designed to mimic the testing dataset. While  Wei, et.al. (2023) only asked GPT-4 to generate similar questions to the GSM8K (Cobbe et al., 2021) dataset, we require GPT-4 to rewrite the question while preserving its original meaning and accuracy. We define a key metric: $\Delta L=L_{test}-L_{ref}$, serving as an indicator of potential test data leakage. A lower $\Delta L$ suggests that the LLM’s lower $L_{test}$ comes from overfitting the test set rather than understanding the questions, indicating potential leakage. Table 6 shows the results of leakage measurement. In addition to the two standard evaluation benchmarks (CEval (Huang et al., 2023) and MMLU (Hendrycks et al., 2021)), we conducted the same experiments on the alpaca dataset (Taori et al., 2023) and the Alpaca-GPT4 dataset (Peng et al., 2023), which is likely used in the pre-training of large models, using its $\Delta L$ as reference. The corpora likely involved in training show a significantly smaller $\Delta L$, whereas the loss for the OpsEval dataset remains at a relatively small positive value. This demonstrates the unbiased nature and non-leakage of the OpsEval test set. The models we used in the leakage test are listed in Appendix B.1.
5.2 Expert alignment of FAE-Score
Table 7: Pearson correlation coefficients between Expert-Evaluation metrics and Automated metrics. Total is the sum of Fluency, Accuracy, and Evidence.
Metric Total Fluency Accuracy Evidence
FAE-Score 0.9175 0.7200 0.9799 0.7962
BLEU-Score 0.6705 0.8253 0.6004 0.4281
ROUGELsum -0.3957 -0.2893 -0.0814 -0.6660
Table 7 shows the correlation coefficients between various automated scoring metrics (ROUGE, BLEU, and FAE-Score) and Expert-Evaluation sub-metrics. The results indicate that ROUGE and BLEU scores often misalign with Expert-Evaluation. This misalignment occurs because LLMs with poor performance may generate keywords that boost ROUGE and BLEU scores, while stronger LLMs might receive lower scores due to different wording from standard answers. In contrast, FAE-Score rankings closely match Expert-Evaluation, particularly with the Accuracy metric. This suggests that FAE-Score is more reliable in assessing the factual accuracy of LLMs’ outputs. Notably, GPT-4’s performance in factual accuracy is reflected in its strong alignment with the Accuracy metric.
6 Limitation and Future Work
Despite the positive contributions of this study, we recognize the following limitations: 1) Topic Distribution Imbalance: There may be an uneven distribution of topic classifications. This issue can be addressed by consciously supplementing with community contributions in future iterations. 2) Agent and RAG Introduction: The inclusion of agents and Retrieval-Augmented Generation (RAG) techniques is constrained by the current large models’ lack of foundational knowledge in operations. Our leaderboard will incorporate more complex tasks once open-source models possess sufficient operational capabilities. 3) Reproducibility of FAE-Score: FAE-Score requires a certain level of instruction-following ability from the judge model, which is why we chose GPT-4 for this work. However, it is best to use open-source models as judge model to ensure reproducibility. In our subsequent work, we will also use open-source, smaller models like Qwen2-72B for evaluation. We believe that any model that ranks higher than Qwen2-72B on general benchmarks, such as SuperGLUE[1], can be considered a suitable judge model. 4) Potential Negative Societal Impact: The use of private domain data from companies necessitates strict adherence to data usage permissions to avoid potential privacy breaches.
7 Conclusion
In this paper, we introduced OpsEval, the first comprehensive Ops benchmark suite designed for evaluating the performance of large language models (LLMs) in IT operations. We established a robust evaluation framework encompassing a wide range of sub-domains and tasks within Ops through rigorous data collection from multiple sources and meticulous preprocessing steps. Our benchmark includes a carefully selected set of 8920 questions, which we have partially released to aid initial evaluations while protecting the integrity of the remaining dataset. It has undergone experiments in data leakage detection, ensuring its reliability. Our observations, supported by quantitative and qualitative results, highlight the need for a balanced approach to selecting fundamental models, considering both performance and robustness. During the QA evaluation, the FAE-Score emerges as a more reliable metric than traditional metrics, suggesting its potential as a replacement for manual labeling in large-scale quantitative evaluations. Our failure rate analysis across 8 tasks and 3 abilities provides researchers with crucial insights and prospects for future breakthroughs.
The identified flexibility within the OpsEval framework presents opportunities for future exploration. This benchmark’s adaptability facilitates the seamless integration of additional fine-grained tasks, providing a foundation for continued research and optimization of LLMs tailored for Ops.
Acknowledgments and Disclosure of Funding
This work is supported by the National Natural Science Foundation of China under Grant 62072264, and the Beijing National Research Center for Information Science and Technology (BNRist) key projects.
Appendix A Details of OpsEval Benchmark
A.1 Information on the companies and experts participating in OpsEval
Table 8: Information of companies collaborating in OpsEval
Organization Domain URL
Bank of Shanghai Financial IT https://www.bosc.cn/zh/
Bizseer Ops service/tool provider https://www.bizseer.com/
ChinaEtek Internet https://www.ce-service.com.cn/
Data Foundation Internet https://www.dfcdata.com.cn/
Guotai Junan Securities https://www.gtja.com/
Huawei Communication https://www.huawei.com/
Lenovo Hybrid Cloud https://www.lenovo.com/
Rizhiyi Log Analysis https://www.rizhiyi.com/
ZTE Communication https://www.zte.com.cn/china/
Zabbix Ops service/tool provider https://www.zabbix.com/
Total 10
Table 8 shows the companies participating in the creation of OpsEval benchmark suite. Their industries include the Internet, telecommunications, cloud computing, finance, and securities, and each company has dispatched at least two experts to participate in the OpsEval work.
A.2 Dependance Filtering Keyword List
question_keywords = [‘the figure’, ‘the scenario’, ‘the previous question’] fail_pred_keywords = [‘unclear’, ‘scenario is not provided’, ‘cannot be determined’, ‘none of the options’, ‘none of the given options’]
A.3 Task Types of Questions
We categorize all questions in OpsEval into 8 tasks. The details of each task are as follows:
  • General Knowledge pertains to foundational concepts and universal practices within the Ops domain.
  • Fault Analysis and Diagnostics focuses on detecting and addressing discrepancies or faults within a network or system, and deducing the primary causes behind those disruptions.
  • Network Configuration revolves around suggesting optimal configurations for network devices like routers, switches, and firewalls to ensure their efficient and secure operations.
  • Software Deployment deals with the dissemination and management of software applications throughout the network or system, verifying their correct installation.
  • Monitoring and Alerts harnesses monitoring tools to supervise network and system efficiency and implements alert mechanisms to notify administrators of emerging issues.
  • Performance Optimization is centered on refining the network and system for peak performance and recognizing potential enhancement areas.
  • Automation Scripts involves the formulation of automation scripts to facilitate processes and decrease manual intervention for administrators.
  • Miscellaneous comprises tasks that do not strictly adhere to the aforementioned classifications or involve a combination of various tasks.
A.4 Ability Levels of Questions
Different questions require different levels of ability to answer. We classify all questions in OpsEval into 3 categories. The details of each ability are as follows:
  1. 1.
    Knowledge Recall: Questions under this category primarily test a model’s capacity to recognize and recall core concepts and foundational knowledge. Such questions are akin to situations where a professional might need to identify a standard procedure or recognize a well-known issue based solely on previous knowledge.
  2. 2.
    Analytical thinking: These questions demand more than mere recall. They necessitate a deeper level of thought, expecting the model to dissect a problem, correlate diverse pieces of information, and derive a coherent conclusion. It mirrors real-world scenarios where professionals troubleshoot complex issues by connecting various dots and leveraging their comprehensive understanding.
  3. 3.
    Practical Application: These questions challenge a model’s ability to apply its foundational knowledge or analytical conclusions to provide actionable recommendations for specific scenarios. It epitomizes situations where professionals are expected to make decisions or suggest solutions based on in-depth analysis and expertise.
Figure 6: Three examples of the processed questions
Figure 6 illustrates examples in our question set, shedding light on our classification methodology.
A.5 Prompt and Formatting of Questions
Figure 7: Three examples of the processed questions
A.6 An Example of Subjective Questions
Figure 8: An example of the saved subjective questions
A saved subjective question in OpsEval is presented in Figure 8, which contains not only the raw question but also its type of task.
Figure 9: An example of building the prompt of subjective questions.
As shown in Figure 9, we combine the task and ability of each question with the question itself as the prompt for LLMs.
A.7 Automated QA generation
During the data collection process, we have experimented automating question-answer generation. We first sampled the QA pairs and manually assessed their accuracy and domain relevance. Later, we used typical manual evaluation examples for few-shot learning, enabling GPT to evaluate QA pairs based on our evaluation criteria automatically. Directly generated question-answers tend to be simple judgment or concept questions rather than reasoning questions that better demonstrate the model’s capabilities and knowledge density. Our goal is to ensure that while the topics of the questions remain relevant to the seed questions, their specific content is distinct from the original questions. By maintaining the overarching framework in the Ops domain, we can expand the number and types of questions, enabling a more comprehensive evaluation of model capabilities. Additionally, we can incorporate external knowledge during the data generation, continually enhancing our ability to evaluate new content.
Appendix B Additional details of experiments
B.1 Detailed Information of LLMs Evaluated
Table 9: Models evaluated in this paper. The “access” column in the table shows whether we have full access to the model weights or can only access them through API.
Model Creator #Parameters Access License
GPT-4/GPT-3.5-turbo OpenAI undisclosed API Proprietary
ERNIE-Bot-4.0 Baidu undisclosed API Proprietary
GLM4/GLM3-turbo Tsinghua Zhipu undisclosed API Proprietary
Meta-LLaMA-3 Meta 8B Weights Llama 3 Community
LLaMA-2 Meta 7/13/70B Weights Llama 2 Community
Qwen-Chat Alibaba Cloud 7/14/72B Weights Qianwen LICENSE
Qwen1.5-Chat Alibaba Cloud 14B Weights Qianwen LICENSE
InternLM2-Chat Shanghai AI Laboratory 7/20B Weights Apache-2.0
DevOps-Model-Chat CodeFuse 14B Weights Apache-2.0
Baichuan2-Chat Baichuan Intelligence 13B Weights Apache-2.0
ChatGLM3 Tsinghua Zhipu 6B Weights Apache-2.0
Mistral Mistral 7B Weights Apache-2.0
Gemma Google 2/7B Weights Gemma license
GPT-4 (OpenAI, 2023a) is a large multimodal model (accepting image and text inputs, emitting text outputs) that, while less capable than humans in many real-world scenarios, exhibits human-level performance on various professional and academic benchmarks. It is recognized as the strongest lanuage model currently. ChatGPT (OpenAI, 2022) is an earlier AI-powered language model developed by OpenAI which is built upon GPT-3.5. We use the GPT-3.5-turbo version in our experiments. LLaMA 2 (Touvron, et.al., 2023) is a second-generation open-source LLM from Meta which is very popular due to its open-source feature. It has the ability to process multiple languages including Chinese. We evaluate three weights (70B, 13B and 7B as shown in 9) of LLaMA 2.
Although LLaMA 2 is able to process Chinese input, it has a small Chinese vocabulary so that its abitilty of understanding and generating Chinese text is limited. As a result, we evaluate some Chinese-oriented LLMs which are published by institutions in China. ERNIE-Bot 4.0 (202, 2024) is the latest self-developed language model released by Baidu. As claimed by Baidu, ERNIE-Bot 4.0 rivals OpenAI’s GPT-4. Qwen (202, 2023) (abbr. Tongyi Qianwen) is a series of LLMs developed by Alibaba Cloud. And Qwen-Chat is a series of large-model-based AI assistant trained with alignment techniques based on the pretrained Qwen. We evaluate three weights (72B, 14B and 7B as shown in 9) of Qwen-Chat. Baichuan2-13B-Chat (Baichuan, 2023) is aligned chat model based on Baichuan2-13B-Base (Baichuan, 2023) which is an open-source LLM published by Baichuan Intelligence. GLM (Du et al., 2022), developed by Tsinghua Knowledge Engineering Group, is a General Language Model pretrained with an autoregressive blank-filling objective and can be finetuned on various natural language understanding and generation tasks. Based on GLM, Zhipu AI released GLM4 (the newest version of GLM model) (Zeng et al., 2022) and GLM3 (the third version of GLM model). For GLM3, we use GLM3-turbo (Zeng et al., 2022) version and ChatGLM3-6B (Zeng et al., 2022) in our experiments. InternLM2-Chat-20B and InternLM2-Chat-7B (InternLM_Team, 2023), recently developed by Shanghai AI Laboratory, are multi-lingual models based on billions of parameters through multi-stage progressive training on over trillions of tokens. Furthermore, we evaluate DevOps-Model-14B-Chat (AI, 2024a), an open source Chinese DevOps oriented models, mainly dedicated to exerting practical value in the field of DevOps.Gemma  (Gemma_Team et al., 2024) is a family of lightweight, state-of-the-art open models based on Gemini technology from Google DeepMind. Trained on up to 6T tokens, Gemma achieves excellent language understanding and reasoning capabilities. We conducted an evaluation of Gemma-2b and Gemma-7b to investigate the effectiveness of Gemma with different weights.
In general, since some models (among them GPT-4, GPT-3.5-turbo, ERNIE-Bot-4.0, GLM4, GLM3-turbo) are not locally available, we evaluate them via API calls. For the remaining models, we perform local inference during evaluation.
Table 10: GPTQ models for LLaMA-2-70B
Model Size #GPTQ Dataset Disc
LLaMA-2-70B 140GB / Raw LLaMA-2-70B model.
LLaMA-2-70B-Int4 35.33GB wikitext 4-bit quantization model.
LLaMA-2-70B-Int3 26.78GB wikitext 3-bit quantization model.
B.2 Prompts
Figure 10: An example of zero-shot evaluation in the CoT setting.Black font represents prompts in English. Purple font represents prompts in Chinese. Red font represents the model’s output in Chinese. Dark red font represents the model’s output in English.
Figure 11: An example of few-shot evaluation in the CoT setting.Black font represents prompts in English. Purple font represents prompts in Chinese. Red font represents the model’s output in Chinese. Dark red font represents the model’s output in English.
For zero-shot evaluation in the CoT setting, we get the answer of LLMs in two rounds. Firstly, by adding a ’Let’s think step by step.’ after the question, LLMs will output its reasoning result. Secondly, we compose the final prompt of the question and the reasoning result in whole as the input of LLMs to get the final answer. An example is shown in Figure 10. For few-shot evaluation in the CoT setting, We make an analysis of each option of the question as a reasoning process, and craft three Q-A examples with CoT reasoning process in answers. An example is shown in Figure 11.
B.3 Compute and Resources Used for Experiments
During our OpEval experiments evaluating different LLMs, we utilize an 8 Nvidia A800-80GB GPU cluster to run inference on models with available weights. For models with API access, we perform inference using CPUs.
B.4 Overview Performance on Different Test Sets
Table 11: LLMs’ overall performance on wired network operations test set
Model English Test Set Chinese Test Set
Zero-shot 3-shot Zero-shot 3-shot
Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC
GPT-4 / / / / / / 88.70 / / / / / / / 86.00 /
Qwen-72B-Chat 70.41 70.50 72.38 72.56 70.32 70.32 70.13 70.22 65.77 65.86 68.13 68.30 69.40 69.40 69.99 70.08
GPT-3.5-turbo 66.60 66.80 69.60 72.00 68.30 68.30 70.90 72.50 58.40 58.60 64.80 67.60 59.20 59.70 65.20 67.40
ERNIE-Bot-4.0 61.15 61.15 70.00 70.00 60.00 60.00 70.00 70.00 67.54 67.54 71.96 71.96 72.00 72.00 78.00 78.00
Qwen1.5-14B-Chat 54.90 34.88 64.09 60.82 52.23 65.55 59.54 47.08 54.04 45.18 62.56 59.12 58.78 61.10 63.43 52.5
Devops-Model-14B-Chat 30.69 30.59 55.77 63.63 63.85 61.96 41.15 44.01 47.59 46.57 52.52 56.01 62.07 60.08 50.59 55.79
Qwen-14B-Chat 43.78 47.81 56.58 59.40 62.09 59.70 49.06 55.88 48.35 48.81 55.35 57.40 58.53 56.12 52.12 54.99
LLaMA-2-13B 41.80 46.50 53.10 58.70 53.30 53.00 56.80 61.00 29.70 31.60 51.60 57.00 39.60 38.90 48.00 50.60
Gemma-7B 25.09 25.09 50.86 50.86 59.12 59.12 50.77 50.77 31.58 31.58 47.59 47.59 34.68 34.68 48.88 48.88
LLaMA-2-70B-Chat 25.29 25.29 57.97 58.06 52.97 52.97 58.55 58.55 38.55 38.55 57.49 57.49 49.09 49.09 48.57 48.57
Internlm2-Chat-20B 56.36 56.36 26.18 26.18 60.48 60.48 45.10 45.10 57.49 57.49 57.14 57.14 59.12 59.12 50.77 50.77
Internlm2-Chat-7B 49.74 49.74 56.19 56.19 48.20 48.20 49.74 49.74 57.49 57.49 57.14 57.14 59.12 59.12 50.77 50.77
LLaMA-2-7B 39.50 40.00 45.40 49.50 48.20 46.80 52.00 55.20 29.80 30.20 50.10 55.60 38.60 40.80 45.60 50.40
Qwen-7B-Chat 45.90 46.00 47.30 50.10 52.10 51.00 48.30 49.80 29.60 29.90 50.60 53.50 50.40 46.90 46.90 47.70
Baichuan2-13B-Chat 37.90 38.30 42.70 46.60 51.90 51.60 44.50 47.45 44.60 45.40 41.60 44.30 45.60 45.70 43.90 46.70
Note: The best accuracy of each language for each LLM is in bold font.
Table 12: LLMs’ overall performance on 5G communication operations test set
Model English Test Set Chinese Test Set
Zero-shot 3-shot Zero-shot 3-shot
Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC
GPT-4 / / 56.30 65.49 / / 59.62 63.54 / / 57.19 62.11 / / 61.55 65.68
Qwen-72B-Chat 53.19 53.19 55.25 55.52 58.13 58.13 58.72 58.99 64.79 64.79 65.79 65.72 70.19 70.19 68.31 68.38
InternLM2-Chat-20B 39.10 39.10 37.70 37.70 47.70 47.70 33.50 33.50 44.60 44.60 47.00 47.00 62.20 62.20 38.30 38.30
Qwen-14B-Chat 33.71 36.25 41.24 42.51 51.19 50.39 57.18 59.18 41.71 41.44 45.58 47.98 53.52 49.92 54.72 58.85
DevOps-Model-14B-Chat 31.04 30.51 42.84 47.37 52.25 49.38 45.90 47.23 41.04 42.70 48.71 53.57 56.85 57.25 51.30 54.29
ERNIE-Bot-4.0 43.66 43.66 51.99 51.99 44.00 44.00 50.00 50.00 45.99 45.99 48.98 48.98 46.00 46.00 54.00 54.00
LLaMA-2-70B 23.64 23.64 39.31 39.31 38.98 39.12 47.90 47.90 24.38 24.38 43.63 43.63 44.65 44.65 48.84 48.84
Mistral-7B 26.91 26.91 30.65 30.65 40.52 40.52 46.84 46.84 1.27 1.27 42.05 42.05 30.72 30.72 46.44 46.44
InternLM2-Chat-7B 36.80 36.80 31.70 31.70 46.30 46.30 36.90 36.90 38.80 38.80 44.60 44.60 46.00 46.00 35.80 35.80
Gemma-7B 23.10 23.10 34.40 34.40 21.40 21.40 33.10 33.10 27.30 27.30 35.40 35.40 17.30 17.30 44.50 44.50
LLaMA-2-13B 15.62 18.32 29.88 34.45 23.16 29.14 37.59 44.3 25.43 27.16 29.17 29.99 36.56 36.15 37.70 39.02
GPT-3.5-turbo 34.92 34.82 38.53 43.50 39.40 39.19 40.93 42.58 36.98 36.83 37.95 39.25 39.17 39.77 41.93 42.15
Qwen-7B-Chat 33.85 33.74 32.45 34.10 32.91 32.70 36.65 36.65 36.27 36.50 33.27 33.51 42.22 40.59 31.28 31.46
ChatGLM3-6B 30.40 30.40 30.70 30.70 26.90 26.90 37.20 37.20 32.60 32.60 35.40 35.40 28.30 28.30 40.90 40.90
Baichuan2-13B-Chat 14.10 15.30 24.10 25.80 32.30 33.10 25.60 27.70 35.64 35.91 30.59 30.52 34.65 35.6 30.21 32.05
LLaMA-2-7B 19.14 21.62 25.70 27.11 21.38 24.85 32.38 34.83 23.57 23.47 27.65 29.26 30.30 30.03 30.98 31.93
Gemma-2B 20.10 20.10 24.20 24.20 31.20 31.20 35.50 35.50 25.60 25.60 28.30 28.30 19.10 19.10 35.50 35.50
Note: The best accuracy of each language for each LLM is in bold font.
Table 13: LLMs’ overall performance on database operations test set
Model English Test Set Chinese Test Set
Zero-shot 3-shot Zero-shot 3-shot
Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC Naive SC CoT CoT+SC
GPT-4 / / 59.02 64.56 / / 58.35 62.58 / / 59.38 65.17 / / 44.06 48.09
InternLM2-Chat-20B / / 59.21 59.21 / / / / / / / / / / / /
ERNIE-Bot-4.0 43.80 43.80 47.14 47.14 46.00 46.00 54.0 54.0 48.56 48.56 50.64 50.64 48.00 48.00 54.0 54.0
Gemma-7B 14.29 14.29 30.99 30.99 2.60 2.60 43.86 43.86 19.32 19.32 53.95 53.95 18.51 18.51 5.20 5.20
Qwen-72B-Chat 47.28 47.48 48.09 48.09 49.70 49.70 43.46 43.66 48.29 48.49 49.50 49.70 49.70 49.70 45.27 44.87
GPT-3.5-turbo 38.63 38.83 40.04 42.05 36.62 37.63 42.66 43.86 36.42 35.81 39.24 43.26 39.84 39.44 27.16 27.77
Qwen-14B-Chat 24.95 28.37 33.00 36.62 27.97 28.37 27.97 24.14 27.57 27.57 32.39 36.02 40.04 35.41 30.38 33.40
DevOps-Model-14B-Chat 25.15 26.96 35.41 38.83 33.20 34.81 27.36 27.36 24.75 22.74 28.37 27.77 36.62 37.02 27.57 26.36
LLaMA-2-70B 19.72 19.72 27.97 27.97 26.56 26.56 32.6 32.6 15.29 15.29 34.81 34.81 26.76 26.76 33.80 33.80
Qwen-7B-Chat 18.91 19.11 22.13 23.94 26.76 25.55 34.81 34.81 18.51 17.71 27.36 28.37 29.78 29.58 33.60 33.60
LLaMA-2-13B 16.10 20.32 23.94 29.58 20.12 22.33 24.35 33.80 23.94 24.35 29.58 31.99 24.55 26.76 21.13 20.72
LLaMA-2-7B 22.13 23.74 23.74 26.56 19.32 20.52 28.77 33.60 20.72 20.72 27.16 27.97 21.53 18.51 18.31 17.91
Mistral-7B 17.10 17.10 26.76 26.76 31.19 31.19 27.97 27.97 0.20 0.20 26.76 26.76 10.26 10.26 32.19 32.19
InternLM2-Chat-7B 27.16 27.16 28.17 28.17 29.98 29.98 30.18 30.18 28.57 28.57 31.79 31.79 30.78 30.78 31.19 31.19
ChatGLM3-6B 20.93 20.93 25.15 25.15 24.75 24.75 29.18 29.18 21.33 21.33 28.97 28.97 21.73 21.73 29.58 29.58
Baichuan2-13B-Chat 17.10 19.11 18.71 22.94 25.96 26.56 20.93 24.55 25.75 25.55 20.12 21.33 27.77 26.76 22.74 24.75
Gemma-2B 16.90 16.90 19.52 19.52 16.10 16.10 24.75 24.75 18.51 18.51 24.95 24.95 21.53 21.53 27.77 27.77
Note: The best accuracy of each language for each LLM is in bold font.
In Table 11, Table 12 and Table 13, we present overview performance of different LLMs on the 3 test sets in OpsEval, including Wired Network Operations, 5G Communication Technology Operations and Database Operations.
B.5 Performance on Different Quantization Models
Figure 12: LLaMA-2-70B’s performance of different quantization parameters. Both zero-shot and few-shot evaluations have been conducted on Wired Network Operations test set under the naive setting.
Figure 12 shows the accuracy of LLaMA-2-70B of different quantization parameters on objective questions, English and Chinese questions respectively. We do both zero-shot and few-shot evaluation with the naive setting.
LLaMA2-70B-Int4 can achieve an accuracy close to LLaMA-2-70B without quantization. Specifically, on English multi-choice questions, the accuracy of the GPTQ model with 4-bit quantization parameters is 3.50% lower in zero-shot evaluation and 0.27% in few-shot evaluation compared to LLaMA-2-70B. As for Chinese questions, the accuracy of LLaMA2-70B-Int4 is 3.67% lower in zero-shot evaluation and 5.18% in few-shot evaluation compared to LLaMA-2-70B. However, LLaMA2-70B-Int3 has a performance degradation that cannot be ignored. On average, the accuracy of LLaMA2-70B-Int3 in English set has a 12.46% degradation compared to LLaMA-2-70B and a 9.30% degradation compared to LLaMA2-70B-Int4.
B.6 Performance on Different Languages
Figure 13: LLMs’ few-shot performance on English/Chinese test set (CoT+SC)
In Figure 13, we compare the few-shot performance of various LLMs under the CoT+SC setting for both English and Chinese questions. Notably, some of the LLMs that have undergone specific training or fine-tuning with Chinese language corpus, such as Chinese-Alpaca-2-13B, Qwen-7B-Chat, and ChatGLM2-6B, still perform better in answering English questions than Chinese ones.
Despite the observed fact that performance tends to be lower for Chinese questions compared to the original English questions, we can still glean valuable insights into the language capabilities of the LLMs. Notably:
  1. 1.
    ChatGLM2-6B experiences the smallest decline in performance when transitioning to Chinese questions. This improvement can be attributed to its substantial exposure to Chinese language data during training rather than simple fine-tuning on top of an existing base model.
  2. 2.
    LLaMA-2-13B exhibits the most significant drop in performance when switching to Chinese questions. This indicates that the shift in language impacts LLMs’ general understanding ability and capacity to extract domain-specific knowledge.
We also observe an interesting phenomenon with the Baichuan-13B-Chat in the 3-shot evaluation with the CoT+SC setting, where its performance in Chinese questions significantly outperforms in English. We examine the LLM’s outputs and analyze a sample question to shed light on this phenomenon in Appendix B.8.4.
B.7 Expert alignment of FAE-Score
Figure 14: Scatter plot and trendline of FAE-Score compared to Expert Evaluation score.
As depicted in Figure 14, the FAE-Score demonstrates a strong positive correlation with Expert Evaluation Score, making it a valuable and effective substitute for automated evaluation.
B.8 Case Study
B.8.1 Failure cases of OpsEval questions
Figure 15: LLMs’ failure cases of OpsEval questions.
Figure 15 illustrates three instances of failure in certain LLMs. The primary reasons for these failures are typically attributed to either a deficiency in domain knowledge or a lack of comprehension of domain knowledge.
B.8.2 Case analysis on QA metrics
Figure 16: Case analysis on QA metrics
Figure 16 illustrates an example where BLEU/ROUGE fail to pick up important keywords and are influenced by the repeated pattern attack, while FAE-Score avoids these two issues.
B.8.3 Case study: Why advanced settings sometimes lack behind
Figure 17: A failed CoT example Generated by InternLM-7B on zero-shot evaluation under the CoT+SC setting in Chinese. Original lines in black are annotated by Chinese version in purple, the dark red and the red text are model’s corresponding output.
Figure 18: A failed 3-shot example Generated by Qwen-7B-Chat on both zero-shot and few-shot evaluations under the naive setting in Chinese.
In certain cases, more advanced evaluation methods surprisingly lead to poorer results. We analyze to understand the potential reasons behind this phenomenon. 1) Some models may respond poorly to the guidance provided by the CoT prompts when required to think step by step, leading to subpar outputs. Figure 17 is one of the examples where CoT failed: the model tested cannot comprehend the idea of thinking step by step. Thus, instead of analyzing each option, it repeated the question and came to its answer directly. Even though the model correctly answered “FTP server” when asked in English, it failed to give the expected option A. This failed case inspires the need for few-shot prompting when applying the CoT method. 2) Few-shot prompts may lead some models to believe that the task involves generating questions rather than answering them, resulting in performance issues. Figure 18 provides an example to the problem mentioned above.
B.8.4 Case study: How Baichuan outperforms in Chinese
Figure 19 shows an example where Baichuan-13B-Chat failed in the English 3-shot CoT+SC setting, with correct English analysis from LLaMA-2-13B and correct Chinese analysis from Baichuan-13B-Chat itself for comparison. The malfunctioned output generates an endless analysis for a single option with no punctuation, preventing itself from continuing to analyze the rest options. This observation suggests that Baichuan-13B-Chat heavily relies on the input language (Chinese in this case) while possessing a foundational knowledge base related to Ops.
Figure 19: A failed English-answering example Generated by Baichuan-13B-Chat on few-shot evaluation under the CoT+SC setting in both English and Chinese.
Appendix C Annotation Guideline for OpsEval Categorization
C.1 Overview
In the OpsEval project, we aim to categorize operational and maintenance tasks within the industry. This categorization process is pivotal for understanding the spectrum of tasks and the required abilities to address them effectively. The process involves two primary steps: automated screening using GPT-4 for initial topic modeling, followed by a manual review process involving domain experts.
C.2 Task Categorization
C.2.1 Objective
To categorize questions into one of eight distinct operational tasks based on industry relevance, task frequency, and significance within operational settings.
C.2.2 Steps
  1. 1.
    Review Initial Categorization: Begin with the insights provided by GPT-4’s topic modeling. Each question has been preliminarily categorized into one or more operational tasks.
  2. 2.
    Understand Task Definitions: Familiarize yourself with the details of the eight distinct tasks outlined in the provided Appendix. Each task has specific criteria and examples to guide your categorization.
  3. 3.
    Assign Tasks: For each question, decide which of the eight tasks it belongs to. A question should be categorized based on its core focus and the operational activity it pertains to.
  4. 4.
    Justification: Briefly justify your choice, especially if a question seems to fit into more than one category. Use the task definitions as a guide to support your decision.
C.2.3 Detailed Task Categorizations
  1. 1.
    General Knowledge: Questions related to foundational concepts and practices in the Ops domain.
  2. 2.
    Fault Analysis and Diagnostics: Questions focusing on identifying and solving discrepancies or faults in systems or networks.
  3. 3.
    Network Configuration: Questions about optimal configurations for network devices to ensure efficient and secure operations.
  4. 4.
    Software Deployment: Questions dealing with the distribution and management of software applications.
  5. 5.
    Monitoring and Alerts: Questions on using monitoring tools to oversee system efficiency and setting up alert mechanisms.
  6. 6.
    Performance Optimization: Questions aimed at enhancing network and system performance.
  7. 7.
    Automation Scripts: Questions involving the creation of scripts to automate processes and reduce manual intervention.
  8. 8.
    Miscellaneous: Questions that do not fit into the above categories or involve elements from multiple categories.
C.2.4 Task Categorization Template
Question ID: Question: [Insert question text here] Assigned Task: Justification: [Provide a brief explanation for the task assignment here]
C.2.5 Example for Task Categorization
Question ID: 001 Question: What steps should be taken to configure a firewall to prevent unauthorized access while allowing legitimate traffic? Assigned Task: Network Configuration Justification: This question specifically asks for optimal configuration strategies for a key network device (firewall) to ensure security and efficient operation, aligning perfectly with the ’Network Configuration’ task.
C.3 Ability Categorization
C.3.1 Objective
To classify questions based on the required cognitive ability to answer them: Knowledge Recall, Analytical Thinking, or Practical Application.
C.3.2 Steps
  1. 1.
    Review Definitions: Read the descriptions of the three abilities in the provided Appendix. Each ability category has distinct characteristics and examples.
  2. 2.
    Evaluate Questions: Assess the cognitive demand of each question. Consider what is primarily required to answer the question effectively: recalling information, analyzing data/situations, or applying knowledge in practical scenarios.
  3. 3.
    Assign Ability Level: Determine the most appropriate ability category for each question. Some questions may seem to require multiple abilities; choose the one that is most critical for addressing the core challenge of the question.
  4. 4.
    Justification: Provide a rationale for your categorization, especially for questions that may not clearly fit into a single category. Refer to the ability definitions to support your categorization.
C.3.3 Detailed Ability Categorizations
  1. 1.
    Knowledge Recall: Requires recognizing and recalling core concepts and foundational knowledge.
  2. 2.
    Analytical Thinking: Demands deeper thought to dissect problems, correlate information, and derive conclusions.
  3. 3.
    Practical Application: Involves applying knowledge or analytical insights to provide actionable recommendations.
C.3.4 Ability Categorization Template
Question ID: Question: [Insert question text here] Assigned Ability: Justification: [Provide a brief explanation for the ability level assignment here]
C.3.5 Example for Ability Categorization
Question ID: 002 Question: How would you optimize the performance of a network experiencing frequent bottlenecks? Assigned Ability: Practical Application Justification: The question requires applying knowledge of network systems and performance optimization techniques to propose specific solutions, hence it falls under ’Practical Application’.
C.4 General Guidelines
  • Consistency: Strive for consistency in your categorization decisions. If similar questions are categorized differently, reassess your choices to ensure they align with the task and ability definitions.
  • Collaboration: When in doubt, discuss challenging questions with fellow experts. Collaboration can help clarify ambiguities and refine the categorization process.
  • Documentation: Keep detailed notes on your decisions, especially for questions that required significant deliberation. This documentation will be valuable for future reference and analysis.
By following these guidelines, you will contribute to a comprehensive and nuanced categorization of operational tasks and required abilities. This effort is crucial for enhancing our understanding of the operational landscape and the diverse skills professionals need to navigate it effectively.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6.
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See Section 6.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No] No, because the reported model performance is significant enough that p-values are not needed to verify the results.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] In Appendix B.3, we present the compute resources used in evaluation.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes]
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] We remove data which contains personally identifiable information or offensive content manually. See Section 3.2.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix C.
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [No] Since this is a community effort, all participants are voluntary and unpaid.